OpenStudy (anonymous):

Expand and simplify. (i represents the imaginary unit) (3-2i)^7

OpenStudy (anonymous):

Please help me!

OpenStudy (kainui):

Simply take \[(3-2i)(3-2i)(3-2i)(3-2i)(3-2i)(3-2i)(3-2i)\]by using FOIL which just means First, Outer, Inner, Last. I'll multiply the first two together (3-2i)(3-2i) and from there you can build on to that. First as in 3*3=9 Outer as in 3*-2i=-6i Inner as in -2i*3=-6i Last as in -2i*-2i=-4 Then add them all together. 9-6i-6i-4=(5-12i) You might be thinking, hey, how come -2i times -2i is -4 and not 4i^2? That's because\[i^2=-1\] Don't forget this. The imaginary unit follows this pattern:\[i^0=1\]\[i^1=i\]\[i^2=-1\]\[i^3=-i\] and then it repeats every 4 powers. So i^4=1 It's a pretty simple pattern to remember as long as you remember that anything to the zeroth power is 1 and anything to the first power is itself. From there, as long as you remember it alternates, you'll know the next two and all others after that by dividing by 4 and finding the remainder.